The weight of a body at rest is 600 N. When the body is in the elevator it has a weight P = 780 N. Determine: a. the acceleration with which the elevator moves. b. The direction of the elevator movement and the type of elevator movement.


1
Expert's answer
2021-01-21T13:02:06-0500

a) Let's apply the Newton's Second Law of Motion:


"P-mg=ma,""a=\dfrac{P-mg}{m},""a=\dfrac{780\ N-600\ N}{\dfrac{600\ N}{9.8\ \dfrac{m}{s^2}}}=2.94\ \dfrac{m}{s^2}."

b) The sign plus means that the elevator moves upward.

Answer:

a) "a=2.94\ \dfrac{m}{s^2}."

b) The sign plus means that the elevator moves upward.